OpenStudy (anonymous):

The voltage V in volts produced by a battery is a linear function of the i in amperes drawn from it .The opposite of the slope of the line repesents the battery 's effective resistance R in ohms .For a certain battery , V=12.0 when i = 10.0 a) What is the effective resistance of the battery ? b) Find the voltage that the battery would producfe when the current is 25.0 amperes.

OpenStudy (anonymous):

key: a)=0.4 ohm b)=2.4 volts

OpenStudy (anonymous):

is that the answer there? a) r=0.4 ohm and v=2.4volts?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok let me fig out something

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

mark.0 did you fig yet?

OpenStudy (anonymous):

my answer is just 2volts not 2.4 volts

OpenStudy (anonymous):

I now a)= (12-8.4)/(1.0-10.0) 3.6/9=0.4 ohm I can't fig b)

OpenStudy (anonymous):

a) = find slope

OpenStudy (anonymous):

25x .4 =10v 12-10=2volts only

OpenStudy (anonymous):

thank, I try many way , I plug 25 for x , 25 for y but not macth, may be book is wrong answer, I will try repost to see if some have difference answer

OpenStudy (anonymous):

ok,,,,good luck cali ....

OpenStudy (anonymous):

t mark o.

OpenStudy (anonymous):

ywc

OpenStudy (anonymous):